1.              If Qs = -20 + 10p, and Qd = 400 - 20p, what is the equilibrium price and quantity? (2 marks)



1
Expert's answer
2021-06-14T13:48:17-0400

"Qd=Qs"

"400-20p=-20+10p"

"400+20=10p+20p"

"420=30p"

"P=14"

"Q=400-20(14)\\=400-280\\=120"